9. A garage offers a breakdown recovery service for motorists that is available every day of the



year. According to their records the number of call-outs they received per day last year were:



Number of call-outs: 0 1 2 3 4


Number of days: 68 103 145 37 12


Required: What is the probability that



A. They receive two call-outs in a day?



B. They receive two or fewer call-outs in a day?



C. They receive one or more call-outs in a day?



D. They receive less than four call-outs in a day?



E. They receive more than two call-outs in a day?

1
Expert's answer
2022-05-24T17:09:53-0400

There are 365 days in the year.

So,

A. "P(=2)=145\/365=0.397"

B. "P(\le2)=P(=0)+P(=1)+P(=2)=68/365+103/365+145/365=0.866"

C. "P(\ge1)=1-P(=0)=1-68/365=0.814"

D. "P(<4)=1-P(=4)=1-12\/365=0.967"

E. "P(>2)=P(=3)+P(=4)=37\/365+12\/365=0.134"